OpenStudy (anonymous):

Write an equation perpendicular to 3x-y=10 through (-6, 1) in standard form.

OpenStudy (azureilai):

3x-y=10 3x=10+y 3x-10=y Your slope right now is 3, in order to find the perpendicular slope, you use the reciprocal. So 3 becomes (1/3) In order to make sure your line passes through the poin, use the point slope formula y-y1=m(x-x1) y-1=(1/3)(x-(-6)) y-1=1/3x+2 y=1/3x+3

OpenStudy (azureilai):

wait its y=-(1/3)x+3. forgot to add signs to the slope.

OpenStudy (anonymous):

Thank you (:

OpenStudy (azureilai):

np